Find a real number, k , such that the line −16x+ky−9=0 has y -intercept −20
Answer by lwsshak3(11628)

Find a real number, k , such that the line −16x+ky−9=0 has y -intercept −20
**
Standard form of equation for a line: y=mx+b, m=slope, b=y-intercept
−16x+ky−9=0
ky=16x+9
y=16x/k+9/k
9/k=-20
k=-9/20